twice the large number is greater than three times the smaller mumber by 19. the sum of the larger number plus one -third of the smaller number is 26 . find the two numbers
Answer by jorel555(1290)

Let m and n be your numbers.
2m-19=3n; 2m-3n=19
m+n/3=26
So:
3m+n=78
9m+3n=234
2m-3n=19
11m=253
m=23
n=9!!!!!!!!!!!!!